OpenStudy (anonymous):

simplify by factoring sqaure root of 40

OpenStudy (anonymous):

\[\sqrt{40}\]

OpenStudy (perl):

40 = 4*10 sqrt(40) = sqrt(4*10) = 2*sqrt(10)

OpenStudy (anonymous):

thanks